### v1

# Scaling Law with Learning Rate Annealing

Abstract
We find that the cross-entropy loss curves of neural language models empirically adhere to a scaling law with learning rate (LR) annealing over training steps ($s$):
$L(s)=L_{0}+A\cdot S_{1}^{-\alpha}-C\cdot S_{2}$
Where $S_{1}$ is forward area and $S_{2}$ is learning rate annealing area. This formulation takes into account two factors: (1) The forward scaling defined as typical scaling law, and (2) the additional loss drop brought by LR annealing. Therefore, this formulation can describe the full loss curve at each step, rather than the single loss point at the end of training. Applying the scaling law with LR annealing and fitting only one or two training curves, we can accurately predict the loss of language model training at any given step and across any learning rate scheduler (LRS). Furthermore, this equation accurately describes the dynamics during training process, and provides a theoretical verification and explanation for numerous experimental findings of previous studies, particularly those focusing on LR schedule and LR annealing. The resulting insights, also serve as a guide for researchers to select critical LRS in advance by prediction using our equation. Most significantly, since all the points in a full training curve follow the equation, we can achieve accurate loss prediction at any given step across any learning rate scheduler, while expending less than 1% of the computational cost required by the chinchilla scaling law to fit language modeling loss. This approach extremely democratizes scaling law fitting and predicting in developing large language models.
1 Introduction
In recent years, the spotlight of academic and industrial interest has been increasingly captured by large language models. The intriguing notion of the scaling law posits that the cross-entropy loss of language models adheres to a power-law pattern as scaling up of model size and data size (Kaplan et al., 2020; Hoffmann et al., 2022). This law serves as a powerful tool, affording researchers the ability to precisely forecast the performance of large language models by fitting the loss at a smaller scale. The potential implications of scaling laws are vast and compelling, sparking a surge of exploration in this burgeoning field (Bahri et al., 2021; Michaud et al., 2023; DeepSeek-AI, 2024).
However, typical scaling laws describe the performance of the model at the end of training, rather than every step during the training process. Essentially, the middle points with different degrees of LR annealing fail to follow typical scaling laws, which do not consider local loss drop brought by LR annealing. Consequently, these laws are unable to fit or predict a full loss curve. Till this work, we do not have an appropriate formulation that accurately describes the dynamics during the training process, which is crucial to deeply understand and improve the training process.
The learning rate schedule (LRS) is a significant variable that greatly influences the loss during the entire training process. Several previous studies have concluded that the rate of decrease in loss accelerates as LR annealing takes place (Loshchilov & Hutter, 2016; Ibrahim et al., 2024; Hu et al., 2024; DeepSeek-AI, 2024). However, these studies generally only provide a qualitative description of how loss changes during LR annealing. Instead, in the present study, we build upon the chinchilla scaling law (Hoffmann et al., 2022) by introducing an additional LR annealing term. This term allows us to accurately predict how loss decreases during LR annealing. Overall, we discover that the loss value at any step of a fixed-size model is determined by two factors: forward area and the degree of LR annealing. Formally, the expectation of loss $L$ at step  [^1] In this paper, we take training steps as data amount in the scaling law equation, which has no essential difference because data amount $=$ training steps $\times$ batch size and batch size is fixed. $s$ of a language model follows:
$L(s)=L_{0}+A\cdot S_{1}^{-\alpha}-C\cdot S_{2}$ (1)
$S_{1}=\sum\limits_{i=1}^{s}\eta_{i}$
$S_{2}=\sum\limits_{i=1}^{s}\sum\limits_{k=1}^{i}(\eta_{k-1}-\eta_{k})\cdot\lambda^{i-k}$
[…]
Figure 1: The definition of $S_{1}$ and $S_{2}$, taking a 20-step cosine learning rate schedule as a toy example. $S_{1}$ (forward area) is the area enclosed by the learning rate curve and the X-axis step, which can be approximately regarded as the total amount of movement of neural network parameters; $S_{2}$ is the decayed annealing area of the learning rate curve (the lighter the color, the greater the degree of decay), that is, the weighted sum of the areas of the small blue squares. Both the increase of $S_{1}$ and $S_{2}$ can help to reduce the loss, but they exist in a delicate balance or trade-off relationship. To minimize the final loss, it’s necessary to discover optimal balance between $S_{1}$ and $S_{2}$.
In Eq. 1, the term $L_{0}+A\cdot S_{1}^{-\alpha}$ is a rectified scaling law, which states that the expected loss decreases as a power-law function of the number of training steps. The term $-C\cdot S_{2}$ is the newly introduced learning rate annealing term. This term represents the additional decrease in loss that occurs as the learning rate is reduced during training. This universal equation can fit the validation loss of any step across any LRS, and then help predict the validation loss of any step across any LRS. As some examples, we fit the equation on the loss curve of constant and cosine LRS in 20K total steps (shown as Fig. 2), then we are able to employ the universal scaling law with LR annealing to predict the full loss curve across various LRS in longer total steps (e.g. 60K) (shown as Fig. 3).
In Sec. 3, we demonstrate the derivation of the scaling law with LR annealing and elucidate the potential theory underpinning our equation. We also extend our equation to a formulation with model size $N$. In Sec. 4, we apply our equation to verify and explain the conclusions drawn from a multitude of previous studies. These conclusions, which previously required substantial computational resources, can now be consolidated in a cost-free manner using our equation. This not only streamlines the process but also brings us closer to understanding the crux of everything about loss drop, LR schedule, and LR annealing (i.e., the art of balance between forward area and annealing area). Moreover, these key insights serve as a guide for researchers to select critical LRS in advance by prediction using our equation. In Sec. 5, we compare our equation with typical chinchilla scaling law (Hoffmann et al., 2022). We show that chinchilla scaling law actually describes special cases (the endpoint of one full loss curve) of our equation, which means that our equation is a generalized form of chinchilla scaling law. Furthermore, we compare the computational costs required by the chinchilla scaling law and our equation, revealing that our equation offers significant savings in computational costs. This greatly democratizes the fitting and prediction of scaling laws.
(a) Learning rate.
(b) Zoomed-out loss fitting.
(c) Zoomed-in loss fitting.
Figure 2: Full loss curve fitting on cosine (20k steps to $\eta_{min}=0$) and constant LRS. The figures omit the warmup in the first 500 steps. After fitting, we get a universal loss equation $L=2.628+0.429\cdot S_{1}^{-0.550}-0.412\cdot S_{2}$.
(a) Full curve prediction of constant LRS.
(b) Full curve prediction of cosine LRS (60K steps to $\eta_{min}=0.1\cdot\eta_{max}$).
(c) Full curve prediction of multi-step cosine LRS (80% + 10% + 10%) (DeepSeek-AI, 2024)
(d) Full curve prediction of WSD LRS (20% cosine annealing to $\eta_{min}=0$(Hu et al., 2024).
(e) Full curve prediction of even strange LRS: periodically repeating linear annealing and linear re-warmup.
Figure 3: Full loss curve prediction (60K steps) by the universal loss curve equation across various LRS. The left, the medium, and the right figures in each row are learning rate curve, zoomed-out loss prediction, zoomed-in loss prediction, respectively. The red rectangle means the zoomed-in zone. The figures omit the warmup in the first 500 steps. The universal loss curve is $L=2.628+0.429\cdot S_{1}^{-0.550}-0.412\cdot S_{2}$, fitted as in Fig. 2. The equation can accurately predict various forms of unseen LRS, and it is highly precise in predicting the trend of loss changes as LR varies. Please note that these are predictive results, which means that none of the points in this figure are involved in the fitting process. The mean prediction errors across various LRS are low to $\sim 0.2\%$.
2 Preliminary
2.1 Scaling Laws
Cross-entropy loss of language models is a good indicator for the performance of almost all downstream tasks (Caballero et al., 2022; Du et al., 2024)Kaplan et al. (2020) empirically discovers a power-law relationship between validation loss and three key factors: model size, dataset size, and the computational resources utilized for training. As an application of scaling law, Hoffmann et al. (2022) trains a compute-optimal large language model named chinchilla by balancing model size and dataset size. Moreover, chinchilla scaling law adopts a simpler and intuitive equation to describe the relationship between the final validation loss ($L$) and the number of parameters ($N$) as well as the amount of data ($D$) as follows:
$L(D,N)=L_{0}+A\cdot D^{-\alpha}+B\cdot N^{-\beta}$ (2)
where $L_{0}$, $A$, $B$, $\alpha$, $\beta$ are all undetermined positive constants. Note that the typical scaling laws fit the loss at the endpoint of training instead of full loss curve, thus a experiment with new data size requires another training launch. Previous works have conducted some preliminary studies on the impact of the learning rate on the scaling laws. For example, OpenAI and chinchilla scaling laws both find that the choice of learning rate schedule does not influence the power-law format (Kaplan et al., 2020; Hoffmann et al., 2022). Also, OpenAI’s experiments suggest that the specific choice of learning rate schedule has minimal impact on the final validation loss, provided that the total summed learning rate is adequately large and the schedule incorporates both a warmup stage and a final annealing stage, reducing the learning rate to nearly zero (Kaplan et al., 2020).
2.2 WSD Learning Rate Scheduler
Hu et al. (2024) proposes a warmup-stable-decay (WSD) LRS including three learning rate stages, which could help get a lower validation loss compared to the typical cosine LRS. The format is like
$WSD(s)=\begin{cases}&\frac{s}{T_{warmup}}\eta_{max},\quad s\leq T_{warmup}\\ &\eta_{max},\quad T_{warmup}<s\leq T_{stable}\\ &f(s-T_{stable})\eta_{max},\quad T_{stable}<s\leq T_{total}\end{cases}$ (3)
Where $0\leq f(s-T_{stable})\leq 1$ is a decreasing function about step $s$, and $\eta_{max}$ is the maximal learning rate. Hägele et al. (2024) consolidates the effectiveness of WSD scheduler by many empirical experiments. Moreover, Hägele et al. (2024) also finds that using 1-sqrt annealing and a moderate annealing ratio (e.g. 20%) can further decrease the final loss.
2.3 Learning Rate Annealing
Learning rate annealing is a common technique used in training neural networks, where the learning rate is reduced over training iterations according to a pre-defined learning rate schedule. This method is often employed to improve the performance and stability of the model during training (Loshchilov & Hutter, 2016). For example, the widely-used cosine LRS could be perceived as a process where the learning rate is cosine curve-like annealing over full steps. Conversely, WSD LRS (Hu et al., 2024) maintains a stable learning rate for the majority of the steps, implementing annealing only in the final (e.g. $10\%\sim 20\%$) of the steps. During the training of LLMs, it has been widely observed that a more pronounced decrease in the learning rate often results in a more precipitous drop in the validation loss (Loshchilov & Hutter, 2016; Ibrahim et al., 2024; DeepSeek-AI, 2024; Hu et al., 2024). However, to the best of our knowledge, all previous studies end with providing a rough and qualitative description of how loss changes during LR annealing, while we provide an equation to formulate the loss changes during LR annealing in this work.
3 Theory
In this section, we elaborate in detail on the intuition, the origin, step-by-step derivation, and the experimental basis of our theory. We then prove the effectiveness of our formula through experiments.
3.1 Similarity between Learning Rate, Gradient Norm, and Loss
Figure 4: The curve shapes of learning rate (top), gradient norm (medium), and validation loss (bottom) hold a quite high similarity across various LRS (labeled as different colors).
As shown in Fig. 4, the first key observation is that the shapes of LR curve, gradient norm curve, and validation loss curve are quite similar across various LRS. This gives a clue that there might exist an implicit connection between learning rate and loss, where gradient norm could be the bridge.
Scaling Laws for Constant LRS.
Constant LRS could be seen as a special LRS, in which any step could be as the endpoint of the LRS. Memorize that chinchilla scaling law (Hoffmann et al., 2022) exactly describes the endpoint of full loss curves. That is to say, the expectation of validation loss of all steps in constant LRS adhere to chinchilla scaling law.
Extra Loss Changes in LR Annealing.
Differently, LR annealing (or re-warmup) brings drastic loss changes in a local range (e.g. Fig. 4), which makes the full loss curve no more follow typical scaling laws. Consider that the LRS without annealing, or constant LRS, follows typical scaling laws. Based on that, we can easily induce that, in a full loss curve, the loss $L$ at step $s$ can be rectified by an added $\eta$-related item, and may follow the format as
$L(s)=\color[rgb]{0,1,1}L_{0}+A\cdot s^{-\alpha}\color[rgb]{1,0,0}-f(\eta)$ (4)
where the blue part is the same as chinchilla scaling law (Hoffmann et al., 2022), while the red part, $f(\eta)$ denotes the extra loss change brought by LR annealing. Recall the curve similarity between learning rate and loss. It could lead to one naive guess: $f(\eta)=C\cdot\eta_{s}$, where $C$ is a positive constant.
Training Discount in Annealing.
The form of Eq. 4 is still imperfect. Note that the gradient $\mathbf{g}$ of parameters decreases along with LR (shown in Fig. 4). Compared with the stages before annealing, the amount of parameter changes (nearly $\eta\cdot|\mathbf{g}|$ of each step) in the neural network optimization process will bring out an almost quadratic faster decline during LR annealing. These clues show that during LR annealing, the loss drop brought by power law (the blue part in Eq. 4) should also decrease. Therefore, we modify to the following format:
$L(s)$ $=\color[rgb]{0,1,1}L_{0}+A\cdot S_{1}^{-\alpha}\color[rgb]{1,0,0}-f(\eta)$ (5)
$S_{1}$ $=\sum\limits_{i=1}^{s}\eta_{i}$
Where we define $S_{1}$ as forward area. $S_{1}$ is exactly the area enclosed by the LRS curve and the X-axis step, which could be approximately regarded as the total amount of neural network parameter changes. Fig. 1 gives a more intuitive visualization for the definition of $S_{1}$.
3.2 LR Annealing Momentum
(a) The different delay steps in the annealing process with different annealing steps (0.1K, 0.5K, 1K and 2K) and the same min learning rate.
(b) The different delay steps in the re-warmup process with different re-warmup steps (0.1K, 0.5K, 1K and 2K) and the same max learning rate.
Figure 5: The delay phenomenon between the learning rate and validation loss. This phenomenon suggests that learning rate annealing (re-warmup) has momentum.
Another key observation is that LR annealing has momentum. For more information about the property of $f(\eta)$, We design one special LRS where the learning rate linearly decreases from $\eta_{max}$ to $\eta_{min}$ then increases. We maintain the slope of the increasing stage (always 5K steps) and change the slope in the decreasing stage. The decreasing steps include 0.1K, 0.5K, 1K, and 2K. Symmetrically, we design another special LRS where the learning rate linearly increases from $\eta_{min}$ to $\eta_{max}$ then decreases. We observe how the validation loss changes with LR increasing or decreasing. The learning rate and loss curves are shown in Fig. 5.
We discover the delay phenomenon between learning rate and validation loss. Firstly, the turning point of validation loss curve is always later than the turning point of learning rate curve in both schedulers. That means that the validation loss still maintains its previous trajectory for a few steps even after the learning rate changes direction. Secondly, the steeper the slope of the learning rate annealing (or re-warmup), the more pronounced the delay between the two turning points. Thirdly, given the same LR slope, the left figure (LR decreasing then increasing) always has a longer delay steps than the right figure (LR decreasing then increasing).
Interestingly, the entire experimental phenomenon is strikingly similar to the physical experiment of a small ball falling along a slope. The larger the angle of the slope, the faster the ball falls. When the ball lands, the accumulated momentum causes the ball to continue sliding for some distance. Based on the delay phenomenon, we pose a hypothesis that $f(\eta)$, the loss drop brought by LR annealing has cumulative historical formation so that the past change of learning rate will affect the following loss curve for a few steps. In summary, learning rate annealing has momentum.
Thus we define $f(\eta)=C\cdot S_{2}$ as follows:
$m_{i}$ $=\lambda\cdot m_{i-1}+(\eta_{i-1}-\eta_{i})$ (6)
$S_{2}$ $=\sum\limits_{i=1}^{s}m_{i}$
$=\sum\limits_{i=1}^{s}\sum\limits_{k=1}^{i}(\eta_{k-1}-\eta_{k})\cdot\lambda^{i-k}$
Where $m_{i}$ is the LR annealing momentum at step $i$, and $\Delta\eta=\eta_{i-1}-\eta_{i}$ denotes the LR annealing amount. $\lambda$ is the decay factor, which signifies the degree to which historical information is retained. We find that $\lambda$ empirically ranges from $0.99$ to $0.999$. In fact, if LR annealing had no momentum, $\lambda=0$ would induce that $f(\eta)=C\cdot\eta_{s}$, which is exactly the naive format mentioned above. Note that $S_{2}$ is not only applicable to LR annealing ($S_{2}>0$), but also to LR re-warmup ($S_{2}<0$). This means that our equation can be also applied in continual pre-training, where LR re-warmup serves as an important factor for better outcomes. More intuitively, the definition of $S_{2}$ can be visualized in Fig. 1, as the weighted sum of the areas of the small blue squares.
3.3 Final Format
Definition.
Given the same training and validation dataset, the same model size, the same training hyper-parameters such as max learning rate $\eta_{max}$ and batch size, the language modeling loss at step $s$ in a full loss curve empirically follow as $L(s)=L_{0}+A\cdot S_{1}^{-\alpha}-C\cdot S_{2}$, where $S_{1}=\sum\limits_{i=1}^{s}\eta_{i}$ and $S_{2}$ are defined as Eq. 6. $L_{0}$, $A$, $C$, $\alpha$ are all undetermined positive constants to be fitted.
It is feasible to apply the formulation to universally describe the loss of each step across diverse learning rate schedulers. This formulation supports validation loss prediction in a more complicated LRS of longer total steps, fitted from a simpler LRS of shorter total steps.
Balance between $S_{1}$ and $S_{2}$.
Notice that $\frac{\partial L}{\partial S_{1}}<0$ and $\frac{\partial L}{\partial S_{2}}<0$ always hold true, which means both the increase of S1 and S2 can help to reduce the loss. However, as shown intuitively in Fig. 1, there exists a delicate balance between $S_{1}$ and $S_{2}$. Once the learning rate begins to anneal and $S_{2}$ starts to increase, the forward area $S_{1}$ of incoming steps starts to diminish. Our equation aptly describes this delicate balance. In Sec. 4, we elaborate in detail on this issue.
3.4 Experiments
LR Warmup.
Different warmup steps can result in different loss curves in training from scratch. During the warmup stage, neural networks are prone to random optimization, resulting in unpredictable outcomes in the very beginning stage (Hestness et al., 2017). One evidence is about the high gradient norm in the LR warmup stage. Fig. 4 shows that at the very beginning of training, the gradient norm initially converges from a relatively high value to a lower one, exhibiting a pattern distinct from the subsequent steps. From another perspective, LR warmup is also necessary. Our pilot experiment (refer to Appendix A) shows that warmup indeed significantly accelerates convergence, a finding also noted by Liu et al. (2020); Kosson et al. (2024). Consequently, in all experiments conducted for this paper, we employ a 500-step learning rate linear warmup to reach the max LR, where $S_{1}$ and $S_{2}$ are computed yet as constant max LR in the warmup stage  [^3] Note LR warmup in training from scratch is different from LR re-warmup in continual training, where we do not regard re-warmup steps as a hyper-parameter and will show how to apply our equation to find optimal re-warmup steps in Sec. 4..
Experimental Setups.
In this paper, we adopt standard experimental setups in large language model pre-training. The training dataset is Fineweb (Penedo et al., 2024) and the validation dataset is RedPajama-CC (Computer, 2023). We train a 594M non-embedding parameters LLAMA architecture-like model (Meta, 2024) from scratch. We use AdamW optimizer (Loshchilov & Hutter, 2017) with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. The weight decay is set as $0.1$ and gradient clip is set as $1.0$. We set maximal learning rate as $2\times 10^{-4}$ and batch size as $4$M tokens. We use the tokenizer of LLAMA-3 (Meta, 2024). We adopt the decay factor of learning rate annealing $\lambda=0.999$ and we discuss the impact of $\lambda$ in Sec. 6.
To validate that our formulation works across different experimental settings, we also conduct our experiments on another experimental setups. Please refer to Appendix B for full experimental setups.
Fitting Details.
Given a learning rate scheduler, we can easily compute out $S_{1}$ and $S_{2}$ of each step in advance. To estimate $(L_{0},A,C,\alpha)$, we adopt a similar fitting method as chinchilla scaling law (Hoffmann et al., 2022). Specifically, we minimize the Huber loss (Huber, 1964) between the predicted and the observed log loss using the L-BFGS algorithm (Nocedal, 1980):
$\min_{L_{0},A,C,\alpha}\quad\sum_{\mathrm{Step~{}}i}\mathrm{Huber}_{\delta}\left(\log\hat{L}(i)-\log L(i)\right)$ (7)
We implement this by the utilization of minimize in scipy library. Huber loss is to enhance to robustness of the fitting results and we set $\delta$ of Huber loss as $10^{-3}$. We mitigate the potential issue of local minima of fitting by choosing the optimal fit from a range of initial conditions. Note that in practice, we can also fit the full loss curves using multiple LRS with a single tuple of $(L_{0},A,C,\alpha)$. In this situation, we sum the Huber losses in Eq. 7 of all fitted LRS.
Fitting and Prediction Results.
We employ the full loss curves from constant and cosine LRS, both containing 20K total steps, as our fitting data points (refer to Fig. 2). We then proceed to predict the full loss curves for various unseen LRS with 60K total steps (refer to Fig. 3). The coefficient of determination ($R^{2}$) in the fitting process is larger than $0.999$, nearly perfect, which underscores the robust capability of our equation to fit loss curves across diverse LRS using a single parameter tuple.
As for prediction shown in Fig. 3, our equation demonstrates broad applicability and robust generalization across five distinct types of LRS, including four previously unseen LRS. This is evidenced by the low mean prediction error, which is approximately 0.2%. Remarkably, our equation is capable to accurately predict losses even in a very complicated LRS with multiple LR re-warmup stages, as illustrated in Fig. 3(e), despite the absence of any LR re-warmup stage in the fitting LRS. We also fit and predict on the curves adopting another experimental setups (shown in Appendix C), and the results are similarly good, which proves that Eq. 1 holds across different experimental setups.
3.5 Extension to Model Size Scaling
Loss Drop in Annealing Stage Scales with Model Size $N$.
First, we research whether the model size $N$ impacts on the amount of loss drop in the annealing stages. We compare the difference of final losses between constant LRS and WSD LRS (10% cosine annealing to $\eta_{min}=0$), thus to get the loss gap brought by LR annealing. We conduct this experiment on different total steps and different model sizes. The experimental results are shown in Fig. 6(a). It suggests that the loss drop brought by LR annealing is scaling with both annealing steps and model size, which means that the annealing area $S_{2}$ in our equation gets larger when model size $N$ becomes large. We suppose there is a simple relationship of $S_{2}\propto N^{\gamma}$ where $\gamma$ is a positive constant to be fitted.
(a) The loss drop brought by LR annealing of different model sizes. The dashed lines represent the trend over steps. It shows that the loss gap brought by LR annealing also scales with data size and model size.
(b) Curve fitting on cosine LRS (60K steps to $\eta_{min}=0.1\cdot\eta_{max}$) of many model sizes using our scaling law extended to model size. By substituting specific values of $N$, the obtained equation simplifies to Eq. 1.
Figure 6: The loss drop brought by LR annealing (on the left) and the $N$-extended full loss curve fitting (on the right). It suggests that $S_{2}\propto N^{\gamma}$ is reasonable from both actual experimental phenomena and curve fitting.
The Format with Model Size.
According to the experiments and analysis above, based on the typical scaling law (Eq. 2), we can extend our proposed equation to model size scaling as follows:
$L(s,N)=L_{0}+A\cdot S_{1}^{-\alpha}+B\cdot N^{-\beta}-C\cdot S_{2}\cdot N^{\gamma}$ (8)
where $N$ denotes the number of non-embedding model parameters (M), and $B$, $\beta$, $\gamma$ are $N$-related undetermined positive constants.
Fitting with Model Size.
We utilize the Eq. 8 to fit full loss curves of different model sizes. The fitting result is shown in Fig. 6(b), where $R^{2}>0.999$ is very close to $1$, suggesting that the loss curves of different model sizes indeed follow our proposed $N$-extended equation. We also fit the curves adopting another experimental setups (shown in Appendix D) and get similar results, proving that Eq. 8 holds across different experimental setups.
4 Takeaways: Experimental Findings Verification and Explanation
Our derived equation describes the training dynamics of language models. In this section, we apply the equation to give theoretical verification and explanation for many existing experimental findings. These key insights also serve as a guide for researchers to select critical LRS in advance by prediction via our equation, with little computational cost. An interesting summary is that
The art of learning rate schedule lies in the delicate balancing act between forward area and annealing area.
4.1 It verifies and explains why loss drops more sharply when LR anneals.
We adopt our equation to help researchers understand why loss drops more sharply when LR anneals, which has been widely observed in many previous studies. We substitute the fitted parameters (see Fig. 2) to our equation as an instance. We draw how the $S_{1}$-item ($A\cdot S_{1}^{-\alpha}$) and the negative $S_{2}$-item ($-C\cdot S_{2}$) impacts the loss along with a WSD scheduler. Fig. 7 suggests that starting from annealing stage, negative $S_{2}$-item has a much more significant impact on the overall loss than $S_{1}$-item, which makes loss drop more sharply compared with the stable LR stage. In conclusion, LR annealing brings out quick increase of the annealing area, resulting in a drastic decrease in validation loss.
Figure 7: How S1-item and negative $S_{2}$-item changes in a WSD scheduler. Gray area means the amount of loss drop brought by $S_{1}$ and $S_{2}$ in annealing stage.
4.2 It verifies and explains when we use cosine LRS, we should set the cosine cycle length $T$ as the total steps $S$, and set min LR as 00 to get the optimal loss.
Figure 8: The predicted loss curves of different cycle length $T$ and min LR in cosine LRS, well aligned with previous studies.
Many papers have found that in LLM pre-training using cosine LRS, setting the cosine cycle length $T$ as the total steps $S$, and setting min LR as nearly 0 (rather than 10% max LR) can lead to the optimal loss (Hoffmann et al., 2022; Hu et al., 2024; Parmar et al., 2024). Actually, the settings above have been a factual standard in LLM pre-training using cosine LRS. We theoretically verify and explain the finding by our equation, presented as Fig. 8. The prediction curve convincingly demonstrates that the loss curve, with the configuration $T=S$ and a min LR of 0, indeed achieves the optimal loss in the end.
Moreover, our equation gives a quite intuitive explanation. $T>S$ leads to an incomplete annealing while $T<S$ leads to a small forward area due to early annealing. Thus, it is optimal to set $T$ as the total steps $S$. Similarly, setting min LR as 00 can make larger annealing amount and thus larger annealing area $S_{2}$, which enables lower final loss.
4.3 It verifies and explains the phenomenon, where constant LRS gets a lower loss than cosine LRS if setting small total steps, and vice versa.
(a) The predicted loss curve of 10K steps.
(b) The predicted loss curve of 100K steps
(c) The predicted final loss of different total steps.
Figure 9: Comparison of constant and cosine in different steps.
In the experiments, we find that if we set small total steps, the final loss of constant LRS could be even lower than cosine LRS, and vice versa. Refer to the ground-truth loss in Fig. 2 (20K steps), and the ground-truth loss in Fig. 3(a)3(b) (60K steps). To validate this phenomenon, we use our equation to draw the prediction loss curve of 10K total steps and 100K total steps in Fig. 9. It shows that our proposed equation can verify well that the better LRS changes over the total steps. Moreover, Fig. 9(c) shows the predicted final loss of different total steps using constant and cosine LRS. It further convincingly suggests that constant LRS indeed gets a lower loss if setting small total steps, but the scaling slope is smaller than cosine LRS’s, resulting in higher loss in more steps.
From a more essential and comprehensive perspective, $|\frac{\partial L}{\partial S_{1}}|$ is a power-law decreasing function while $|\frac{\partial L}{\partial S_{2}}|$ is stable over training steps. In the early stages, $|\frac{\partial L}{\partial S_{1}}|$ is large when $S_{1}$ is small, thus increasing $S_{1}$ by maintaining large LR (e.g. constant LRS) in the early stages can greatly help reduce the loss. That is, $S_{1}$ plays a dominant role over $S_{2}$. In the later stages, $|\frac{\partial L}{\partial S_{1}}|$ is much smaller when $S_{1}$ becomes large, thus increasing $S_{1}$ in the later stages does not significantly help reduce the loss. That is, $S_{2}$ plays a dominant role over $S_{1}$. At this stage, It is time to start LR annealing to increase $S_{2}$. Interestingly, this perspective aligns directly with the idea of WSD LRS (Hu et al., 2024): In the early stages, the neural network is exploring globally and it is a suitable time to use a larger LR; In the later stages, the neural network is exploring locally and it is a suitable time to use a smaller LR. We will delve further into WSD LRS in the following subsections.
4.4 It verifies and explains WSD and multi-step cosine LRS have matched or even lower loss than cosine LRS.
(a) Learning rate curves of three types of LRS.
(b) $S_{1}$-item and negative $S_{2}$-item of different LRS.
Figure 10: The comparison between $S_{1}$-item and negative $S_{2}$-item in different LRS.
Recently, it has been demonstrated that WSD LRS (Hu et al., 2024) and multi-step cosine LRS (DeepSeek-AI, 2024) yield a lower loss compared to the typical cosine LRS. We also prove this by the experiments (refer to the ground-truth loss in Fig. 3(b)3(c)3(d)). We validate and elucidate this finding using our proposed scaling law with LR annealing. In Fig. 10, we depict the learning rate (on the left) and the predicted loss drop (on the right) for different LRS using our derived equation. The figures suggest that for multi-step cosine or WSD LRS, the negative $S_{2}$-item ($-C\cdot S_{2}$) is slightly higher than that of cosine LRS. However, the $S_{1}$-item ($A\cdot S_{1}^{-\alpha}$) is significantly lower than that in cosine LRS. More simply to say, both WSD LRS and multi-step cosine LRS aim to employ a strategy that marginally reduces $S_{2}$ but substantially increases $S_{1}$, leading to an overall decrease in validation loss.
4.5 It verifies and explains that a moderate WSD annealing ratio could get the lowest loss.
In the case of WSD learning rate schedule, it is crucial to ascertain the optimal annealing ratio for training steps. Prior research by Hägele et al. (2024) has established the existence of an optimal annealing ratio within the WSD scheduler. As illustrated in experimental results of their study, excessively high or low annealing ratios lead to sub-optimal model performance. This phenomenon can be further elucidated through our proposed equation.
Theoretically, a high annealing ratio results in a significant reduction of the forward training area $S_{1}$, while only marginally increasing the annealing area $S_{2}$. Conversely, an excessively low annealing ratio leads to under-annealing, characterized by a diminutive $S_{2}$ and consequently get a high final loss. Our scaling law function builds a trade-off relationship between the forward area $S_{1}$ and annealing area $S_{2}$ about the annealing ratio.
As depicted in Fig. 11, we utilize the fitted scaling law function to draw the final loss across various annealing ratios and total training steps. The prediction results present parabola-like curves, and align well with the actual experimental outcomes by previous works. It suggests that a moderate WSD annealing ratio is optimal which could make a moderate $S_{1}$, and then maximize $S_{1}$ and $S_{2}$ from a global view, thereby minimizing the overall validation loss. Moreover, our equation directly help predict an optimal annealing ratio for different total steps without experiments, which saves a lot of resources.
(a) The relationship between the predicted final loss and the ratio of annealing steps under the condition of different total steps.
(b) The relationship between predicted final loss and the forward area $S_{1}$ of different total steps. Different points denote different annealing ratios.
Figure 11: Illustration of the predicted loss in relation to the ratio of annealing steps and the forward area in WSD LRS (cosine annealing), presenting parabola-like curves, with a distinct optimal loss.
4.6 It verifies and explains that the optimal annealing function in WSD LRS depends on the annealing ratio.
(a) The predicted loss curve of cosine and 1-sqrt annealing method of different annealing ratio.
(b) The true loss curve of different annealing ratios with cosine and 1-sqrt annealing methods.
Figure 12: The predicted (left) and true loss (right) of cosine and 1-sqrt annealing method at different annealing ratios. Experimental results (right), aligned with our prediction (left), refute the previous finding “the order and results of different annealing hold across settings” (Hägele et al., 2024).
In the context of the WSD LRS, the selection of the annealing method in the annealing stage is also pivotal to optimize the training process. Hägele et al. (2024) conclude that the 1-sqrt annealing (refer to Appendix E for 1-sqrt function and curve) yields a lower final loss compared to the other annealing methods (e.g. cosine). They claim that the conclusion holds true across different annealing ratios.
However, as we predict using our equation (Fig. 12(a)), it indicates that 1-sqrt annealing does get a lower loss than cosine annealing in 10% and 20% annealing ratios, but performs much worse than cosine annealing in 50% annealing ratio.
To verify whether the predictions from our equation are accurate, we conduct experiments by training models using different annealing methods and ratios within a fixed 50K total steps. As illustrated in Fig. 12(b), at a 10% or 20% annealing ratio, the 1-sqrt method outperforms the cosine method, whereas at a 50% annealing ratio, the latter method exhibits a lower final loss. The true experimental results align quite well with our prediction, which also overturns some of the conclusions made by previous works. We conclude that the optimal annealing function in WSD LRS depends on the annealing ratio.
Our scaling law function provides an explanatory framework for these observations. We draw the LR curves of 1-sqrt and cosine annealing in Appendix E. At 10% annealing ratio, although the forward area $S_{1}$ of the cosine method is slightly larger than that of the 1-sqrt method, the larger annealing area $S_{2}$ of the 1-sqrt method plays a more critical role in reducing the overall final loss. However, as the annealing ratio increases, the difference of $S_{1}$ between two LRS gradually becomes larger and larger, till breaking the delicate balance between $S_{1}$ and $S_{2}$ at 50% annealing ratio, resulting in a lower final loss for the cosine method. This relationship underscores the importance of carefully selecting the annealing strategy to optimize model training outcomes within the WSD scheduler. Still, our equation can help predict a better annealing method without experiments, which saves a lot of resources.
4.7 It verifies and explains that in continual pre-training, the higher max learning rate to re-warmup, the higher the initial peak loss will be, and then the more sharply it will decrease.
(a) The predicted validation loss with different re-warmup max LR in the continual pre-training process. All the re-warmup steps are 500 steps.
(b) The predicted validation loss with different re-warmup max learning rate in the continual pre-training process.
Figure 13: The predicted validation loss with different re-warmup max learning rate and re-warmup steps in the continual pre-training process. The LRS of continual pre-training is cosine ($T$=100K) and the min learning rate is 0.
In continual pre-training (CPT), the learning rate scheduler is usually set as re-warmup to a new max LR at the beginning. By many experiments, Gupta et al. (2023) concludes that the higher max learning rate to re-warmup, the higher the initial peak loss will be, and then the more sharply it will decrease.
According to our scaling law function  [^4] Strictly speaking, continual pre-training process often include LR re-warmup as well as data distribution shift. Here we primarily research on the condition where there is no distribution shift between two training stages. The conclusions transfer across most cases because the loss change brought by LR re-warmup is significantly larger than the loss change brought by data distribution shift (Gupta et al., 2023; Ibrahim et al., 2024)., in the re-warmup process, the annealing area $S_{2}$ will reduce to a negative value ($S_{2}<0$) and thus the validation loss increases. The higher max LR in re-warmup, the annealing area $S_{2}$ becomes more negative and thus there would be a higher peak loss. But still, higher max LR could make the forward area $S_{1}$ grow faster and the loss decreases more sharply after re-warmup. We use the fitted equation to predict the continual pre-training process with different max LR as shown in Fig. 13(a). The predicted loss curves reproduce a quite similar phenomenon with previous works (Gupta et al., 2023).
There is a more profound strategy using our equation in CPT. As shown in Fig. 13(a), after ensuring total steps during CPT, we can apply our equation to predict a better max LR and scheduler to get the lowest final loss without experiments, which saves a lot of resources.
4.8 It verifies and explains that in continual pre-training, the steps of re-warmup have little impact on the final loss.
Meanwhile, how many steps to re-warmup is another important issue in the continual pre-training.  Gupta et al. (2023) find that the longer re-warmup steps could smooth the transition of loss curve but the number of re-warmup steps does not significantly influence the final validation loss. We use the fitted equation to predicted the continual pre-training dynamics with different re-warmup steps. The results, shown in Fig. 13(b), present a good alignment with previous works (Gupta et al., 2023).
Based on our theory, given the fixed max LR, when the re-warmup steps are longer, the annealing area decreases more slowly and the loss curve rises more smoothly, but both final $S_{1}$ and $S_{2}$ are quite stable across different re-warmup steps. First, the annealing area $S_{2}$ of different re-warmup steps are very close due to the same max LR and the same min LR. Besides, though different re-warmup steps bring in temporary distinct losses, re-warmup only cover a small percentage compared with all training steps. Thus, the forward area $S_{1}$ is also close across different re-warmup steps, resulting in the close overall loss across different steps of re-warmup.
5 Comparison with Chinchilla Scaling Law
5.1 Reduction to Chinchilla Scaling Law
Our scaling law function can predict the full loss curve across any given learning rate scheduler. In this section, we show that our equation has no contradiction with typical scaling law, and it is a generalized form of the chinchilla scaling law (Hoffmann et al., 2022). That is to say, all the final loss points of different total training steps following our equation should also follow a power-law relationship. We prove this by dividing into two conditions: (1) constant LRS, and (2) other LRS.
Constant LRS.
In the case of constant learning rate scheduler, the annealing area $S_{2}$ is always zero and the forward area $S_{1}=\eta_{max}\cdot s$, where $s$ is the step, and $\eta_{max}$ is the maximal learning rate which is constant over steps. Thus, the whole train loss curves becomes:
$L(s)=L_{0}+(A\cdot\eta_{max}^{-\alpha})\cdot s^{-\alpha}=L_{0}+A^{\prime}\cdot s^{-\alpha}$ (9)
which can be easily reduced to the chinchilla scaling law equation.
Other LRS.
For other learning rate schedulers, we adopt a method based on statistics to show that our scaling law function can be reduced to the chinchilla scaling law. Specifically, we check whether chinchilla scaling law fits well the endpoints of loss curves predicted and generated by our scaling law. The parameter tuple of our equation is $(L_{0},A,C,\alpha)$. We then randomly sample 1000 sets of parameter tuples in some uniform distributions: $L_{0}\sim U(1,3)$, $A\sim U(0.3,0.5)$, $C\sim U(0.2,0.6)$, $\alpha\sim U(-0.6,-0.4)$. Each parameter tuple could be seen as the fitting result of a distinct set of experimental setups  [^5] It’s worth noting that some of these sampled parameter tuples might not be reasonable or likely to happen in real-world scenarios, but we choose to keep them nonetheless.. (e.g. dataset, batch size, model size, etc.). For each generated parameter tuple, we apply our equation to predict the final loss of different total training steps range from 5K steps to 60K steps. We conduct the prediction on two LRS including cosine and WSD (10% annealing ratio). The predicted final loss points are used to fit the chinchilla equation through minimizing the Huber loss. Fitting examples are shown in Fig. 14.
(a) The predicted loss of different total steps with cosine LRS and the fitted chinchilla curve.
(b) The predicted loss of different total steps with WSD LRS and the fitted chinchilla curve.
Figure 14: Chinchilla scaling law fits well the validation loss endpoints predicted by our formulation, taking cosine LRS (on the left) and WSD LRS (on the right) as examples.
LR Scheduler $\text{mean}(R^{2})$uparrow $\text{std}(R^{2})$downarrow Huber loss $\downarrow$
Cosine 0.972 0.056 0.00017
WSD 0.979 0.053 0.00013
Table 1: Mean and standard deviation of $R^{2}$, as well as the mean Huber loss of $1000$ randomly generated parameter tuples.
We calculate the mean and standard deviation of $R^{2}$ and Huber loss among 1000 random parameter tuples. As shown in Table 1, the random selected parameter tuples have low Huber loss and high $R^{2}$ value. The experimental results demonstrate that chinchilla scaling law fits well on the data predicted and generated by our scaling law function. That is, our scaling law with LR annealing can be reduced to chinchilla scaling law.
5.2 Scaling Law Fitting Democratization
We extremely democratize the fitting and prediction of scaling law by greatly enlarging its scope of applicability. Adopting our scaling law with LR annealing, we can collect thousands of fitting points in one full loss curve during training, while typical scaling law can only collect one endpoint from a full loss curve.
For a more direct comparison with chinchilla scaling law (Hoffmann et al., 2022), we suppose a scenario where it requires 100 fitting points to get the parameters of scaling laws. We assume the distance between each point as $K$. We compute the required training steps using different approaches as follows:
  • Adopting chinchilla scaling law, typical cosine LRS requires total steps of at least $1K+2K+3K+\cdots+100K=5050K$.
  • Adopting chinchilla scaling law, WSD LRS (notating annealing ratio as $r$) requires total steps of at least $(1K+2K+3K+\cdots+100K)r+100K(1-r)=(100+4950r)K$.
  • Adopting our scaling law, all we need is only one training curve with moderate total steps (and the number of fitting points is far more than $100$), such as one curve with $50K$ steps  [^6] The empirical rule that more fitting points always achieve better fitting results always holds true. Our equation can also use more points and LRS for fitting, such as $30K$ constant + $70K$ cosine. Nevertheless, we can collect far more fitting points than the typical scaling law with significantly fewer training steps..
We present a comparison of the computational costs associated with different laws and LRS in Table 2. Our findings indicate that our proposed equation enables us to fit loss with less than 1% of the computational cost required by the chinchilla scaling law. Furthermore, our scaling law with LR annealing, can be universally applicable to other LRS by fitting any given LRS (except constant), thereby conserving additional computational resources. This approach extremely democratizes the process of fitting and predicting scaling laws in LLM pre-training, paving the way for a more environmentally friendly and carbon-efficient methodology.
Equation LRS Computational cost Applicable to other LRS?
Chinchilla Cosine 100% No
Chinchilla WSD (20% annealing) 21.6% No
Chinchilla WSD (10% annealing) 11.8% No
Ours Any (except constant) $<$1.0% Yes
Table 2: The comparison of computational cost for fitting, assuming $100$ fitting points as an example.
6 Discussion
6.1 The impact of Decay Factor $\lambda$
(a) Learning rate curves.
(b) The loss curve of $\lambda=0.99$.
(c) The loss curve of $\lambda=0.999$.
Figure 15: The comparison of fitting effect of different decay factor $\lambda$.
Figure 16: The predicted loss in different annealing ratios of WSD LRS for $\lambda=0.99$ and $\lambda=0.999$.
The decay factor $\lambda$ in our equation plays a crucial role to indicate the information retaining degree in LR annealing. We set $\lambda$ as $0.999$ in our all experiments. We explore the difference from another decay factor $\lambda=0.99$. We fit and get different equations for different $\lambda$. We compare them (1) on the predicted loss curves for 1-square and 1-sqrt annealing methods, and (2) on the predicted loss curves in different annealing ratios of WSD LRS (cosine annealing).
The results, illustrated in Fig. 15 and 16, reveal several key insights into the impact of decay factor:
Delay Steps.
A larger decay factor results in longer delay steps. Comparing Fig. 15(b) and Fig. 15(c), $\lambda=0.999$ introduces a more obvious delay phenomenon, which is consistent across both the 1-square and 1-sqrt annealing methods. The root reason is simple: larger $\lambda$ can retain more LR historical momentum, causing longer delay steps after LR finish annealing.
Optimal Annealing Ratio.
a larger decay factor tends to favor a higher annealing ratio. As shown in Fig. 16, The optimal annealing ratio of $\lambda=0.999$ is larger than that of $\lambda=0.99$. Meanwhile, due to the similar reason, $\lambda=0.999$ favors 1-sqrt annealing method while $\lambda=0.99$ favors 1-square annealing method, as shown in Fig. 15.
Balance Point between $S_{1}$ and $S_{2}$.
More essentially, the selection of $\lambda$ decides the balance point of $S_{1}$ and $S_{2}$. For example, $\lambda=0.999$ means that, LR annealing only retain the information of previous approximately $\frac{1}{1-\lambda}=1000$ steps, which can be seen as the window size of LR annealing momentum. The window size could be very close to the optimal annealing steps. After reaching window size, $S_{2}$ increases very slowly, with the cost of large decrease of $S_{1}$.
The analyses above underscore the sensitivity of our scaling law function to the choice of decay factor and highlights the importance of selecting a decay factor that aligns closely with empirical data to ensure the accuracy of predictions. We recommend that the future developers try different $\lambda$ for their own setups  [^7] Actually, $\lambda$ can be fitted as a parameter, instead of a hyper-parameter requiring manual tuning. We regard $\lambda$ as a hyper-parameter because $\lambda=0.999$ performs well in our most experiments. Besides, fitting with $\lambda$ could bring in additional time complexity due to the recomputation of $S_{2}$..
6.2 Possible Root Reasons of Delay Phenomenon in Learning Rate Annealing
In Sec. 3, we discover the delay phenomenon, which proves that LR annealing has momentum. We discuss possible root reasons of the phenomenon in this section.
Adam Optimizer? No.
We notice that Adam optimizer (Kingma & Ba, 2015) also has the first-order momentum decay factor $\beta_{1}$ and the second-order momentum decay factor $\beta_{2}$, which presents the possible connection to the the delay phenomenon.
We keep $\beta_{1}=0.9$, and conduct delay experiments on different $\beta_{2}\in\{0.95,0.99,0.999\}$ (default: $0.95$) of AdamW optimizer (Loshchilov & Hutter, 2017) to observe whether larger $\beta_{2}$ causes a more longer delay steps. The learning rate and ground-true loss curve are shown in Fig. 17(a). It suggests that the ground-truth loss curves of different $\beta_{2}$ almost coincide with each other, and their delay steps are also the same. Therefore, we believe that Adam optimizer has little to do with the delay phenomenon, despite its momentum form seeming very related to our experiments. Speaking of which, we even once tried to mimic the form of Adam Optimizer to describe LR annealing momentum, attempting to discover a connection between them, but the fitting results were a mess.
(a) The comparison of true loss curve with setting different $\beta_{2}$ of Adam optimizer.
(b) The comparison of delay steps of different re-warmup steps (and thus different $S_{1}$).
Figure 17: The possible root reason analysis (left: Adam optimizer, right: $S_{1}$) of delay phenomenon.
Forward Area $S_{1}$? Not Really.
No matter how LR changes, $S_{1}$ is always increasing over steps, resulting in consistently reducing the validation loss brought from $S_{1}$. Therefore, the forward area, $S_{1}$ would lengthen delay steps in LR annealing then re-warmup, but would shorten delay steps in LR re-warmup then annealing. The delay phenomenon is indeed related to $S_{2}$.
But still, $S_{1}$ is not all the reasons of delay phenomenon. We prove this by Fig. 5(b), which suggests that even though in LR re-warmup then annealing, the delay phenomenon, while not that pronounced, still exists. Moreover, we conduct delay experiments by adjusting the slope of LR after tuning point of LR. As shown in Fig. 17(b), We find that more smooth slope of LR re-warmup, with smaller $S_{1}$, but still causes longer delay steps. Therefore, we conclude that $S_{1}$ indeed influences the specific delay length, but is not the root reason.
Other Possible Reasons?
The delay phenomenon could be intuitive in some cases. For example, suppose that learning rate decreases directly from 2e-4 to 2e-5 in one step, and maintains 2e-5. In this case, although the loss would decrease to a lower value but the parameter changes in one step is too small in neural networks. Given a sudden low LR, neural networks still require some steps to gradually optimize to a local minimum, incurring delay phenomenon. But still, analysis above still ends with a rough description, and we have not figured out the root reasons and look forward to completing this part in future work.
6.3 Other Possible Scaling Law Formats with LR annealing
Adding a LR-weighted Coefficient to $S_{2}$?
Imagine that when LR anneals to nearly 00, the neural network’s parameters almost do not change and the validation loss should not change, either. However, as defined in our equation, Eq. 1, $S_{2}$ still has historical momentum even if LR is nearly 00, making the loss continue to decrease and misalign with observed training dynamics.
To cover this corner case, we try a revision to our equation and add a LR-weighted coefficient to $S_{2}$. Specifically, we adjust $S_{2}$ to more approach 00 when $\eta$ is close to 00, counteracting the original formulation’s tendency to overestimate loss reduction when $\eta\approx 0$.
The revised equation for the annealing area $S_{2}$ in our scaling law function is as follows:
$m_{i}$ $=\lambda\cdot m_{i-1}+(\eta_{k-1}-\eta_{k})$ (10)
$=\sum\limits_{k=1}^{i}(\eta_{k-1}-\eta_{k})\cdot\lambda^{i-k}$
$S_{2}$ $=\sum\limits_{i=1}^{s}m_{i}\color[rgb]{1,0,0}\cdot\eta_{i}^{\epsilon}$
Where the red part is the added LR-weighted coefficient and $\epsilon$ is a undetermined positive constant. $\epsilon$ could be very small in practice.
We have tried the revised function to fit data. We find that the fitting results are quite similar and $\epsilon$ is very close to 0, showing little use in practical effect. Hence, we adopt the original format in our experiments  [^8] We still recommend future developers to try this format if possible..
$L\propto S_{2}^{\zeta}$ rather than $L\propto S_{2}$?
(a) LR of cosine and constant.
(b) The fitting curve of $L\propto S_{2}$
(c) The fitting curve of $L\propto S_{2}^{\zeta}$.
Figure 18: The comparison of fitting effect between $L\propto S_{2}^{\zeta}$ with $L\propto S_{2}$.
Actually, all we know is that $L$ and $S_{2}$ have a positive correlation. Thus $L\propto S_{2}^{\zeta}$ rather than $L\propto S_{2}$ might be a more reasonable assumption. That is, our equation would be changed to $L(s)=L_{0}+A\cdot S_{1}^{-\alpha}-C\cdot S_{2}^{{\color[rgb]{1,0,0}\zeta}}$. Theoretically, the introduction of $\zeta$ as an additional fitting parameter is expected to provide a more nuanced control over how changes in the learning rate annealing affect validation loss, potentially leading to improve the accuracy of our equation.
However, the empirical results, as depicted in Fig. 18, demonstrate that the fitting improvement with the inclusion of $\zeta$ is quite marginal when compared to the version without this parameter. This slight enhancement does not justify the additional complexity introduced by managing negative values of $S_{2}$. Furthermore, the empirical observation that $\zeta$ tends to converge close to $1$ (e.g. $1.125$ in Fig. 18(c)) reinforces the notion that the original formulation of the function, without the power term $\zeta$, is adequately robust. This finding suggests that the direct influence of the learning rate annealing area, as initially modeled, sufficiently captures the essential dynamics without the need for this additional complexity  [^9] Another additional complexity lies in that $S_{2}^{\zeta}$ becomes incalculable when $S_{2}<0$ in LR re-warmup..
7 Future Works
7.1 Tuning Scaling Law Format
We have tested a variety of equation formats to enhance the accuracy of the entire training process. As a result, the final equation format, as presented in Eq. 1, proves to be optimal so far across a range of scenarios. And the equation has achieved a level of practicality that enables the prediction of future loss when scaling training steps and model sizes. We will persist in refining the equation format to forecast the loss with the highest possible accuracy.
7.2 Extension to Post-training
In this work, we research primarily on the scope of pre-training of LLM. We also show how to apply our equation to guide the LR re-warmup strategy in continual pre-training. We will continue researching on how to extend our equation to post-training, which might include data distribution shift, data mixture, model alignment, and specific downstream evaluations.
8 Conclusion
In conclusion, we discover that the loss curves of neural language models empirically conform to a scaling law with learning rate annealing over training steps ($s$), which is expressed as $L(s)=L_{0}+A\cdot S_{1}^{-\alpha}-C\cdot S_{2}$. This equation points out the art of balance between forward area and annealing area of learning rate schedulers, offering theoretical verification and explanation for numerous experimental findings of prior studies. We present the underlying theory of our equation including curve similarity and LR annealing momentum. We extend our equation to the format with model size $N$. Moreover, we prove that our equation is a generalized form of typical scaling laws, further showing the superiority of our equation.
Compared to typical scaling laws which predict the endpoint of one full loss curve, our equation can accurately forecast the loss of language model training at any given step and across any learning rate scheduler. As for required computation resources, our equation expends less than 1% of the computational cost required by typical scaling laws to fit language modeling loss. Our proposed approach greatly democratizes the fitting and prediction of scaling laws in the development of large language models.
Appendix A Impact of Warmup Steps
We conduct experiments on the imapct of learning rate warmup steps. As shown in Fig. 19, we find that 500 warmup steps can get the lowest validation loss compared to 100 or no warmup steps. The experimental results also guide us to choose 500 warmup steps in our experiments of this work.
Figure 19: The comparison of the true loss curve of different warmup steps. We experiment on cosine LRS with 20K total steps.
Appendix B Experimental Setups
In this work, we use multiple sets of experimental setups, in order to validate that our equation can work across different experimental setups. For clarification, we present the experimental setup list as shown in Table 3.
In our most experiments (including our takeaway sections), we use the main setting A. In our fitting and prediction experiment, we repeat experiments and consolidate our conclusion using another setting, setting B, and the experimental results are shown in Fig. 21. In our $N$-extended scaling law fitting experiment, we repeat experiments and consolidate our conclusion using another setting, setting C, and the experimental results are shown in Fig. 22.
Setups Setting A (mainly) Setting B Setting C
Model Size (Non-embedding) $594$M $293$M multiple
Train Dataset Fineweb Finweb Mixture-train*
Val Dataset RedPajama-CC RedPajama-CC Mixture-valid*
Total Steps $60$K $120$K $143$K
Maximal Learning Rate $2\times 10^{-4}$ $2\times 10^{-4}$ $1.381\times 10^{-3}$
Warmup Steps $500$ $100$ $500$
Batch Size (tokens) $4$M $2$M $4$M
Tokenizer LLAMA-3’s LLAMA-3’s Extended LLAMA-2’s
$\beta_{1}$,$\beta_{2}$ in Adam Optimizer $0.9$, 0.90.90.90.95 $0.9$, $0.95$ $0.9$, $0.95$
Weight Decay $0.1$ $0.1$ $0.1$
Gradient Clip $1.0$ $1.0$ $1.0$
Table 3: Experimental settings adopted in this work. * denotes pre-training multilingual dataset including mixture of sources such as common crawls, books, arxiv, code, etc. “Extended LLAMA-2’s” is extended from LLAMA-2’s tokenizer  (Meta, 2023) by adding vocabulary.
Appendix C Our Scaling Law on Another Experiments Setups
Fig. 2 and Fig. 3 show that our equation can work very well on our main experimental setup. For proving that our scaling law with LR annealing can apply to different (but given) experimental settings, we change the setting from $A$ to $B$ (refer to Table 3) and observe whether our equation can still work or not. The fitting results are shown in Fig. 20. The prediction results are shown in Fig. 21. The results suggest that our scaling law with LR annealing can still work well across different experimental setups.
(a) Learning rate.
(b) Zoomed-out loss fitting.
(c) Zoomed-in loss fitting.
Figure 20: Full loss curve fitting on cosine (30K steps to $\eta_{min}=0$) and constant LRS. The figures omit the warmup in the first 100 steps. After fitting, we get a universal loss equation $L=2.761+0.517\cdot S_{1}^{-0.491}-0.458\cdot S_{2}$. Refer to setting B in Table 3 for experimental setups.
(a) Full curve prediction of constant LRS.
(b) Full curve prediction of WSD LRS (90K total steps; 10 % cosine annealing to $\eta_{min}=0$).
(c) Full curve prediction of WSD LRS (110K total steps; 10 % cosine annealing to $\eta_{min}=0$).
Figure 21: Full loss curve prediction (120K steps) by the universal loss curve equation across various LRS. The left, the medium, and the right figures in each row are learning rate curve, zoomed-out loss prediction, and zoomed-in loss prediction, respectively. The red rectangle means the zoomed-in zone. The figures omit the warmup in the first 100 steps. The universal loss curve is $L=2.761+0.517\cdot S_{1}^{-0.491}-0.458\cdot S_{2}$, fitted as in Fig. 3(b). The equation can accurately predict various forms of unseen LRS, and it is highly precise in predicting the trend of loss changes as LR varies. Please note that these are predictive results, which means that none of the points in this figure (except constant LRS) are involved in the fitting process. The mean prediction errors across various LRS are low to $\sim 0.2\%$. Refer to setting B in Table 3 for experimental setups.
Appendix D Our $N$-extended Scaling Law on Another Experiments Setups
Fig. 6(b) show that our $N$-extended equation can work very well on our main experimental setup. Similarly, for proving that our $N$-extendec caling law can apply to different (but given) experimental settings, we change the setting from $A$ to $C$ (refer to Table 3) and observe whether our equation can still work or not. The fitting results are shown in Fig. 22. The results suggest that our $N$-extended scaling law with LR annealing can still work well across different experimental setups.
Figure 22: Curve fitting on cosine LRS (143K steps to $\eta_{min}=0$) of many model sizes using our scaling law extended to model size $N$. Refer to setting C in Table 3 for experimental setups.
(a) LR curve of WSD (20% 1-sqrt/cosine annealing).
(b) LR curve of WSD (50% 1-sqrt/cosine annealing).
Figure 23: The learning rate curves of 20% (left) and 50% (right) annealing ratio in WSD LRS, with cosine and 1-sqrt annealing method.
Appendix E 1-sqrt and 1-square Annealing
The 1-sqrt annealing is proposed by Hägele et al. (2024), whose learning rate at step $s$ in annealing stage is defined as:
$f(s)=1-\sqrt{\frac{s-T_{\text{stable }}}{T_{\text{total}}-T_{\text{stable }}}}$ (11)
$\eta_{s}=\eta_{min}+f(s)\cdot(\eta_{max}-\eta_{min})$
We draw the learning rate curve of WSD (20% and 50% 1-sqrt annealing) in Fig. 23, compared with cosine annealing. Other than 1-sqrt annealing, Hägele et al. (2024) also mentions 1-square annealing method defined as:
$f(s)=1-\left({\frac{s-T_{\text{stable }}}{T_{\text{total}}-T_{\text{stable }}}}\right)^{2}$ (12)
$\eta_{s}=\eta_{min}+f(s)\cdot(\eta_{max}-\eta_{min})$